N2 gas is effused through a pinhole in 5.0 sec. Determine the time it will take an 

equivalent amount of CO2 to effuse under the same conditions.


1
Expert's answer
2020-08-02T15:49:34-0400

Solution.

"(\frac{M(CO2)}{M(N2)})^{0.5} = \frac{t(CO2)}{t(N2)}"

"(\frac{44.0095}{28.0134})^{0.5} = \frac{x}{5.0}"

x = 6.3 sec.

Answer:

t(CO2) = 6.3 sec.